OpenStudy (anonymous):

7x - 14 = 4x + help me please explain properly

hero (hero):

7x - 14 = 4x + what?

OpenStudy (anonymous):

7

hero (hero):

7x - 14 = 4x + 7 7x - 4x = 14 + 7 3x = 21 x = 21/3 x = 7

OpenStudy (anonymous):

i get it thanks

hero (hero):

:D

hero (hero):

I figured you would

OpenStudy (anonymous):

hehe you a math teacher or sumink ?

hero (hero):

Math Major

OpenStudy (anonymous):

haha lol

OpenStudy (anonymous):

7x - 14 = 4x + 7 7x - 4x = 7 + 14 3x = 21 x = 21/3 = 7

hero (hero):

Hey look! That's the exact same thing I did above. How do I know you didn't copy and paste?

OpenStudy (anonymous):

lol